Write the series 2 + 9 + 28 + 65 + ⋯ in Sigma notation. 


1
Expert's answer
2022-04-12T06:48:59-0400

"2=1^3+1\\9=2^3+1\\28=3^3+1\\65=4^3+1\\\\2+9+28+65+...=\sum_{n=1}^{\infty}{\left( n^3+1 \right)}"